# Overcoming Data Scarcity in Generative Language Modelling for Low-Resource Languages: A Systematic Review

### v1

@@ 4.2.6. Adaptive Learning
In terms of adaptive learning, two papers opted to extend the training of preexisting models by exposing them to Turkish text data (Toraman, 2024; Acikgoz et al., 2024). This interesting approach takes advantage of cross-lingual information through transfer learning and subsequently yields promising results. One of the papers highlights that continued training with raw text and that extending the vocabulary size did not have an impact on performance (Toraman, 2024). The other paper interestingly highlights that resuming the training process on multilingual models can trigger catastrophic forgetting of their primary language knowledge (Acikgoz et al., 2024). While exploiting transfer learning could improve performance for generative language models, the paper indicates that large models like Gemma7B still perform better without fine-tuning. This could indicate that better results can be obtained by collating a larger more representative dataset and building models with more parameters. Word alignments have also been added for instruction tuning to enhance the performance in translation tasks for LRLs, indicating that instruction fine-tuning could be an effective measure of overcoming data scarcity in LRL translation tasks (Mao and Yu, 2024).
@@ 4.2.7. Family of Low-Resource Languages
A variety of papers target building generative language models for families of related languages (Niyogi and Bhattacharya, 2024; Cahyawijaya et al., 2021; Wongso et al., 2023; Downey et al., 2024; Li et al., 2022a). One paper successfully builds multilingual language models for the Uralic language family (Downey et al., 2024). Although the study indicates that LRLs benefit from being trained as part of a wider Uralic multilingual language model, even the more resourced languages such as Hungarian and Russian also benefit. They suggest it is infeasible to model some LRLs such as Komi and Moksha without exploiting multilingualism and transfer learning. This chimes with the results of IndoNLG (Cahyawijaya et al., 2021), as their model achieves similar results to another large multilingual model (mBART-large), even though they use  20% of the number of parameters. One of the previously mentioned studies that experiments with 250 high and LRLs (Chang et al., 2023) suggests that building multilingual models can be compared to augmenting the dataset for a monolingual model by 33%. It could be both beneficial and insightful to investigate the impact of training multilingual models for closely related languages. It would be interesting to observe whether a model trained on a family of LRLs could be compared to an even larger increase in data. These results suggest that building generative language models for a family of LRLs could not only yield strong performance but also reduce the need to train separate models for each language, making the approach more cost-effective and environmentally sustainable.
@@ 4.2.8. Mass Translation
Mass translation has been used to generate parallel data by employing services like Google Translate API and GPT 3.5/4 (Shafayat et al., 2024). As mentioned previously, mass translation can be used for boosting monolingual data and also for augmenting parallel data. While this process has been executed to create QA systems for Bengali (Shafayat et al., 2024), the results are highly dependent on the translation quality of these translation services. Although similar methods are used to bolster LRL datasets for translation, this process could generate inaccurate text that could impede QA and other generative tasks such as dialogue generation. Another paper uses translation tools to impute missing data in parallel datasets (Haque et al., 2021). While this could be an effective way of dealing with missing values in a dataset, it could also corrupt the semantic information in text data which would result in inaccurate information being generated for tasks such as QA.
@@ 4.2.9. Miscellaneous
One unique method mentioned within this review pertains to extending the vocabulary size of existing generative models that have been predominantly trained on English textual data (Toraman, 2024). While this is an interesting approach, the paper finds that adapting existing models by extending the vocabulary does not have a significant impact on performance. However, it highlights that continual training can improve perplexity. This could indicate that additional training outweighs the vocabulary size when training generative language models for LRLs. In addition, one paper applies the concept of code-mixing as a method of augmenting text data (Li et al., 2024b). While this approach is previously discussed within the scope of being a multilingual approach, code-mixing should also be referenced with its capacity to capture more community-centric language variations. Although code-mixing can be used to build models that generate text in multiple languages, code-mixing in this way is not necessarily modelling each of these languages, but it is rather modelling the intersection of them. The paper suggests that this approach can create synthetic data with a Code-Mixing Index comparable with existing authentic datasets. This could be significant for building generative language models for LRLs that closely relate to another language. Given that English and Irish terms are often used interchangeably in Ireland (Laoire, 2016), a code-mixing approach could subsequently build a generative language model that more accurately captures the colloquial language of that community. Not only has code-mixing been shown to be effective in addressing data scarcity when building generative language models, but it can also enhance the representation of communities that naturally interweave multiple languages into their vernacular.
@@
Given the prevalence of data augmentation in computer vision tasks, monolingual text augmentation techniques should be explored in further detail. Furthermore, this category encapsulates a discussion on the algorithms and processes presented in the results section for RQ3, that could be used to synthesise new training data for LRLs.
@@ 4.3.1. Paraphrasing
Paraphrasing was the most referenced data augmentation technique in our review set of papers. Several papers employ existing or additional models to paraphrase authentic text data and subsequently synthesise new sentences that preserve the semantic information of the original data (Maimaiti et al., 2021; Sun et al., 2019; Guo et al., 2021a; Gao et al., 2020). One paper creates paraphrased text by using randomly generated text sequences as input to a model trained on Tibetan QA pairs to generate new synthetic QA pairs (Sun et al., 2019). The study subsequently employs BERT to apply some word-level grammatical fixes to the output text. While the results of their generation model indicate that some text inputs can produce coherent QA pairs, these results can also ignore the grammar characteristics of the target language. Although this approach of utilising a language model to paraphrase monolingual text data can produce grammatical errors, this could be compared to the distortions that can occur with back-translation. Therefore, it could be insightful to enlist paraphrasing tools to bolster text data in the context of machine translation tasks. In contrast, this potentially erroneous and nonsensical output could have a detrimental effect on the training of models for other generation tasks like dialogue generation and abstractive text summarisation.
@@
A variety of studies in this review produce paraphrased text by enacting different types of word replacement strategies (Haque et al., 2021; Gao et al., 2019; Shi et al., 2022; Wang et al., 2018; Maimaiti et al., 2021). Another approach refers to creating numerous duplications of original sentences and applying random replacements on the duplicated sentences to produce new paraphrased sentences (Guo et al., 2021a). Much like the word-level grammatical checks used to improve the quality of generated QA pairs (Sun et al., 2019), performing word-level augmentations by replacing terms with random or context-informed words could also distort the original interpretation of the sentence. One of these approaches tackles the shortcomings of word-level replacements by utilising POS tag-based augmentations (Maimaiti et al., 2021). This is an interesting process as choosing specific POS tags and replacing them could reduce the risk of grammatical errors produced through synthesising data. If a language has word classes that do not change based on agreement with other words in a text, a contextualised word replacement based on POS tags could produce more grammatically correct augmented sentences. For example, adverbs do not affect agreement in English sentences, therefore they could be replaced with semantically similar terms without corrupting the grammar of the sentence. However, it is worth highlighting that this approach is highly dependent on the quality of available POS tagging tools and models for a given LRL. Not only does the quality of POS-informed replacements hinge on the availability of POS tagging tools, the efficacy of the replacement strategy also directly impacts the quality of the paraphrased text. Although POS-based augmentations could reduce the risk of synthesising grammatically incorrect text, the replacement strategy, neural-based or not, could corrupt or dilute the original meaning of the text being augmented. While this POS-informed approach is used to improve NMT for LRLs, this data augmentation method could also bolster general language generation models, particularly for LRLs that consist of word classes which do not affect the grammar of other terms in a sentence.
@@ 4.3.2. Grammatical Transformation
There are a couple of papers that perform reordering on monolingual sentences as an approach for synthesising new data from existing text data (Wu and Wang, 2023; Marie and Fujita, 2020). While these methods can be viewed as a form of paraphrasing, they explicitly perform a degree of restructuring. These approaches involve duplicating existing sentences and reconstructing them as a means to augment the entire dataset. One paper uses a reverse transcription grammar model to generate diverse sentences for given input sentences (Wu and Wang, 2023). While this method can produce grammatically correct text, the reordering process could corrupt the original meaning of the sentence. Therefore, it is possible that this method could be beneficial for NMT tasks in the same way that back-translation improves translation performance. In terms of generating text for dialogue generation or abstract summarisation, training text data containing many distortions of factual content could produce poor models. However, this process of reordering a sentence could be beneficial for LRLs that have word order flexibility. Another approach relates to synthesising new text data by introducing noise to reformat input text data to resemble user-generated text (Marie and Fujita, 2020). This is achieved by applying various stylistic changes such as introducing spelling errors, slang and informal expressions. While this approach is applied to improving NMT systems for user-generated content, there could be applications of this method for building grammar alteration models. Models could be built for adapting or translating from one sentence structure to another within the same language. This would still follow the sequence-to-sequence nature of NMT but subsequently produce various augmentations of an original input sentence. Not only could producing such synthesised sentences improve NMT, but this form of text data augmentation could also empower other generative tasks like dialogue generation and QA tasks for LRLs.
@@ 4.3.3. Data Enrichment
The approach of enriching existing data by appending additional information to it is quite a straightforward approach that relates to adding the titles of books to each example sentence for building a translation model from Historical Japanese to Contemporary Japanese (Usui and Komiya, 2023). While the impact of this data enrichment technique is not explored in depth, more research is required to determine if adding metadata to text documents could help with text generation tasks. It could be worth exploring the impact of including other data such as the access date, the year, the authors and the title of the training data. Performing this process could help to ground the generative language model through the inclusion of relevant information for each text example.
@@ 4.4. RQ4: Which Sources Are Publishing in the Area of Low-Resource Language Modelling?
The results section for RQ4 indicates that there is a wide variety of sources publishing in the area of LRL modelling. Given the dominance of ACL publications in this review spanning journals, workshops and conferences alike, it is clear that it has a central influence on the dissemination of literature relating to modelling LRLs. The high frequency of publications from specific events such as EMNLP and LREC could suggest that these are key venues for LRL research. The prevalence of EMNLP papers could be attributed to the empirical nature of language modelling, and the frequency of LREC papers could be linked to the language resource element of building language models for LRLs. While ACM is the publisher of fewer events in comparison with other organisations such as ACL and IEEE, the Transactions on Asian and Low-Resource Language Information Processing journal is the most frequently sourced publication in this review. This could highlight that this journal is highly influential in terms of LRL modelling. It is also worth mentioning the prevalence of arXiv preprints in this review. While arXiv papers are not necessarily associated with traditional academic journals or undergo formal peer review, their prevalence could suggest that the field of LRL modelling is a rapidly evolving research topic. Given that arXiv papers consist of 19% of the review dataset, they could introduce limitations relating to the quality of research in the review process. However, we mitigate this effect by enlisting the quality assessment question relating to the rankings of conferences and journals. On the other hand, IEEE has more of a multi-disciplinary influence as it is the source of eight different publications relating to various fields within ICT and engineering. This could highlight that there are potential applications for LRL modelling in a variety of domains. Considering that Springer is the source of five papers in this review, this could indicate that there is demand for journals that relate to specialised fields to include literature that tackles LRL modelling. Single contributions and participation from other publishers such as Frontiers, IJCAI, and De Gruyter could indicate that the research area of LRL modelling could be rising in prevalence. Ultimately, there is a broad range of sources that span eight different publishers in the field of LRL modelling.
@@ 4.5. RQ5: What Architectures Are Being Used to Produce Low-Resource Language Models?
The prevalence of transformer-based models in this review reflects the transformative nature of the architecture for generative language modelling. Given the rise in popularity, adoption and performance of transformer-powered services such as ChatGPT, it is unsurprising that it dominates as the most frequently implemented model architecture in this review. Figure 13 indicates that transformer-based models could be the most suitable architecture for modelling LRLs. RNNs, and more specifically LSTMs are also somewhat popular architectures for modelling LRLs in this review. This could reflect the evolution of the state-of-the-art architecture over time as RNN/LSTM-based approaches would have been the gold standard before the transformer was introduced (Mienye et al., 2024). While GAN, GRU and SMT-based methods have been used by studies in this review, they are not as popular for modelling generative models for LRLs. Their unpopularity could suggest that they are not the optimal architectures for LRL generation. Interestingly, GPT-2 is the most popular model across all of the studies seen in this review. Given that GPT-2 is an open-source predecessor to the widely popular GPT-3 and GPT-4 models by OpenAI, its popularity could indicate that it is suitable for LRL modelling. The size of GPT-2 small could make it appealing to LRL researchers, as it does not require as much training time and computational power to build it as other prevalent models such as the Llama3 family of models (Radford et al., 2019; Dubey et al., 2024). XLM and sequence-to-sequence models such as mT5 and T5 are also frequently used for LRL modelling. Their popularity could be attributed to the considerable number of models in this review that relate to translation, as the task is naturally a sequence-to-sequence task. Ultimately, there is a wide variety of models being used to build generative models for LRLs, although transformer-based approaches are the most frequently used for this task.
@@
While some text data augmentation techniques are effective at improving the performance of different NLG models, it is worth considering how such methods compare across different NLG tasks. Translation as a generative language task can be improved with a number of augmentation approaches such as back-translation (Pham et al., 2021; Hong et al., 2024; Berckmann and Hiziroglu, 2020; Yirmibeşoğlu and Güngör, 2023; Nissanka et al., 2020; Dandapat and Federmann, 2018; Pham Van and Le Thanh, 2022), code-mixing (Bhowmick et al., 2023) and paraphrasing (Marie and Fujita, 2020; Guo et al., 2021a). It is unlikely that back-translation would be an effective method for bolstering the amount of training data for building models for other LRL NLG tasks such as summarisation or dialogue generation. Using back-translation for data synthesis could introduce inconsistencies at a syntactic and semantic level in the training corpus. Given that synthesising code-mixed data has yielded performance increases for translation tasks, it could be insightful to apply this approach to languages that are often interwoven with each other. For example, modern Gaeilge usage could be more accurately modelled if it was trained on a corpus code-mixed to include English, given the strong relationship between both languages (Laoire, 2016). Therefore, it could be interesting to investigate the impact of code-mixing on other NLG tasks. While paraphrasing has produced performance improvements for translation tasks, it has also yielded improvements for other tasks such as question-answering corpus generation and dialogue generation (Sun et al., 2019; Gao et al., 2020; Wu and Wang, 2023). This approach of using paraphrasing models to bolster training data is a promising approach that hinges on the availability of enough data or existing paraphrasing models. More research and experimentation is required to assess the impact of large-scale paraphrasing as a means to augment training data for building NLG models for LRLs. Training language models on a family of related languages has been quite effective at bridging the data gap between LRLs. Given that this language family-based approach can reduce the number of model parameters (Cahyawijaya et al., 2021) and that multilingual modelling can be compared to a 33% increase in training data for an LRL (Chang et al., 2023), it is clear that this is an effective text data augmentation method. It would be interesting to apply this language family-based approach to various groups of related LRLs to further validate this observation.

### v2

@@ 4.2.6. Adaptive Learning
Adaptive learning has been leveraged by several studies to model LRLs. While two papers extended the training of pre-trained models by exposing them to additional Turkish data (Toraman, 2024; Acikgoz et al., 2024), this approach is prone to significant challenges. As explored in the monolingual modelling section, further training an existing pre-trained model allows it to benefit from transfer knowledge but it still needs to train on a significant amount of text data for models with billions of parameters. In addition to needing a large amount of data to avoid overfitting, the pre-trained models must support the tokens required to represent a given LRL. However, some evidence shows that resuming training with new raw text and extending the vocabulary size does not have an impact on performance (Toraman, 2024). Further research could be done to confirm these findings for other LRLs.
One study observed that resuming training can trigger catastrophic forgetting of the primary language knowledge of multilingual models (Acikgoz et al., 2024). This could suggest that the data used to continue training needs to be representative of the existing split of languages learned by the model. On the other hand, if monolingual modelling is the objective then catastrophic forgetting could be ideal as a multilingual model adapts to a target language during training. The same study indicates that large models like Gemma7B still perform better without fine-tuning. This could suggest that collating a larger more representative dataset and building models with more parameters could be more beneficial for modelling multiple LRLs at once.
An open challenge would be to explore ways of extending models to account for new changes in language. Without ongoing updates, models will not capture emerging language, such as slang or cultural references (e.g., a model trained before 2024 would not understand the cultural context behind the phrase “brat summer”, popularised by British singer Charli XCX (Guardian, 2024)). While adaptive learning can address data scarcity in the short term, future work should investigate how it can also support long-term model evolution, particularly for low-resource languages.
@@ 4.2.7. Family of Low-Resource Languages
Although families of LRLs are used less than both multilingual and monolingual modelling, the approach is heavily linked to multilingual modelling. The difference in popularity may reflect that not all multilingual modelling approaches are explored using related languages. However, the findings suggesting that multilingual modelling can yield performance gains equivalent to a 33% increase in training data, should be further investigated in the context of related LRLs (Chang et al., 2023). This improvement could be enhanced by selecting similar or maximally related languages for multilingual modelling. This argument draws on the evidence that shows more resourced languages as well as LRLs can benefit from being modelled with related languages (Downey et al., 2024; Cahyawijaya et al., 2021). The reduction in required parameters through modelling families of related languages (Cahyawijaya et al., 2021) should also be investigated further as reducing model sizes could lower both the cost and environmental impact of LRL modelling.
We agree that it could be infeasible to model some extremely LRLs (such as Komi or Moksha) without exploiting multilingualism and transfer learning across languages (Downey et al., 2024). Therefore, this necessitates further research focusing on pooling dataset resources across related linguistic communities. Future work should also investigate modelling families of LRLs as a way of reducing model sizes. This branch of research could promote environmentally friendly methods of solving the data scarcity issue for LRLs.
@@ 4.2.8. Mass Translation
Using tools like Google Translate API and GPT 3.5/4 as a form of mass translation is one of the least popular methods used for overcoming LRL data scarcity. Its lack of popularity is likely linked to the same pitfalls facing other previously discussed translation-based approaches, like back-translation. The major pitfall relates to the constraints of existing translation tools. Augmenting a monolingual dataset by leveraging a translation service to bulk translate from a source language to a target language limits the performance of subsequent models. In other words, any translation errors would propagate through to any models trained on the synthetic data. While one study uses this approach to build QA systems for Bengali (Shafayat et al., 2024), these results would be constrained by the translation service’s capacity to translate Bengali text data. For a task like QA where factual information could be queried, losing semantic content during translation could produce corrupted information.
Although mass translation is not a widely favoured approach to addressing data scarcity, exploring the use of translated data to train generative models could still serve as a useful baseline for comparison with models trained on authentic text. Future work could explore translation methods that focus on preserving semantic content during translation. Ensuring translation systems do not distort factual information may result in a more useful approach for augmenting LRL data.
@@ 4.2.9. Miscellaneous
While code-mixing could be considered a form of multilingual model, it is also tagged with the miscellaneous category as it differs in its output. Code-mixing could model languages or dialects that use terms from multiple languages interchangeably. Given that Irish speakers often inject English into their Irish phrases and vice versa (Laoire, 2016), future work could investigate using code-mixing to augment data for representing bilingual communities.
Additionally, extending the vocabulary of existing pre-trained models is related to adaptive learning but it is also tagged with the miscellaneous category. Given that pre-training from scratch can be costly and time-consuming, adapting existing models that may not have tokenisation support for LRL scripts is a challenge. However, some evidence highlights that extending the vocabulary size of a model along with continuing training on new text might not have an impact on performance (Toraman, 2024). Although this study did not find any evidence of performance changes, further research could investigate the impact of the quantity of new data when extending the vocabulary size on model performance.
@@
Given the prevalence and success of data augmentation in computer vision tasks, monolingual text augmentation techniques should be explored in further detail. Therefore, this section contains a discussion on the algorithms and processes presented in the results section for RQ3, that could be used to synthesise new training data for LRLs, in order of method frequency.
@@ 4.3.1. Paraphrasing
Paraphrasing techniques can be grouped into two primary categories: (1) word or sentence-level alterations, and (2) generation-powered paraphrasing..
@@
Word-level replacement strategies vary from study to study, however, they all share common risks. Much like the word-level grammatical checks used to improve the quality of generated QA pairs (Sun et al., 2019), performing word-level augmentations by replacing terms with random or context-informed words could also distort the original semantic content of the sentence. The use of POS tag-based augmentations (Maimaiti et al., 2021) is an alternative process as choosing specific POS tags and replacing them could reduce the risk of grammatical errors produced through synthesising data. If a language has word classes that do not change based on agreement with other words in a text, a contextualised word replacement based on POS tags could produce more grammatically correct augmented sentences than random replacement. For example, adverbs tend not to affect agreement in English sentences, therefore they could be replaced with semantically similar terms without corrupting the grammar of the sentence. However, this approach is highly dependent on the quality of available POS tagging tools and models for a given LRL. Not only does the quality of POS-informed replacements hinge on the availability of POS tagging tools, the efficacy of the replacement strategy also directly impacts the quality of the paraphrased text. POS-based augmentations could reduce the risk of synthesising grammatically incorrect text, but the replacement strategy, neural-based or not, could corrupt or dilute the original meaning of the text being augmented. While this POS-informed approach is used to improve NMT for LRLs (Maimaiti et al., 2021), this data augmentation method could also improve general language generation models, particularly for LRLs that consist of word classes which do not affect the grammar of other terms in a sentence. Future work could explore POS-informed replacements as a method for increasing language diversity across different generation tasks.
Alternatively, generation-powered paraphrasing has shown limited results in generating new synthetic QA pairs (Sun et al., 2019). However, the study’s limited success in generating coherent text, even with BERT-powered grammar correction, is undermined by grammar characteristics being ignored in the target language. Although this approach of utilising a language model to paraphrase monolingual text data can produce grammatical errors, this could be compared to the distortions that can occur with back-translation. Therefore, it could be insightful to adopt paraphrasing tools to bolster text data in the context of machine translation tasks. In contrast, this potentially erroneous and nonsensical output could have a detrimental effect on the training of models for other generation tasks like dialogue generation and abstractive text summarisation where grammatical cohesion is expected. Although this generation method is prone to errors, LRL translation tasks could benefit from it as these errors are similar to the kind of grammatical errors that back-translation can produce.
@@ 4.3.2. Grammatical Transformation
This discussion section is split between providing a meta-analysis of sentence restructuring and introducing noise to augment monolingual data as forms of grammatical transformations.
The grammatical transformations category involves studies using a degree of sentence restructuring to produce synthetic sentence-level data. The use of a reverse transcription grammar model to generate diverse sentences for given input sentences can produce grammatically correct text (Wu and Wang, 2023), however, the reordering process could corrupt the original meaning of the sentence. Therefore, it is possible that this method could be beneficial for NMT tasks in the same way that back-translation can improve translation performance. In terms of generating text for dialogue generation or abstract summarisation, training text data containing many distortions of semantic content could produce poor models. However, this process of reordering a sentence could be beneficial for LRLs that have word order flexibility. Furthermore, this grammatical transformation method could be used to produce synthetic, parallel data for dialects that vary in word order from a standard form of a language. Future work could investigate the quality of synthetic text generated by reordering sentences to reflect dialect-specific syntax or word order. This would be beneficial as it could augment data for translation tasks between dialects.
Alternatively, the success of applying noise to improve user-generated sequence-to-sequence translation performance (Marie and Fujita, 2020) should be explored for other generative modelling tasks. While this approach applies various stylistic changes such as introducing spelling errors, slang and informal expressions, the approach could be used for building grammar correction models. That is, this sequence-to-sequence approach could be used to fix grammatical agreement errors in sentences, where sentences with incorrect agreement could be mapped to sentences with correct agreement. Furthermore, datasets could be built for adapting or translating from one dialectal sentence structure to another within the same language. A dataset with a sentence in one dialect mapped to another could train a model for generating synthetic, parallel data between two language dialects. This would still follow the sequence-to-sequence nature of NMT but subsequently produce various augmentations of an original input sentence. This form of text data augmentation could effectively empower other generative tasks like dialogue generation and QA tasks to ensure grammatical quality and dialectal diversity for LRLs.
@@ 4.3.3. Data Enrichment
While enriching existing data by appending additional information is a relatively straightforward approach, it remains underexplored in the literature (Usui and Komiya, 2023). Given that the impact of this data enrichment technique is not explored, more research is required to determine if adding metadata to text documents could help with text generation tasks. Including other data such as the access date, the year, the authors and the title of the training data could help to ground the generative language model through the inclusion of relevant information for each text example. However, this data enrichment approach does not introduce new semantic content; instead, it enhances the depth of the existing text data. Future work should empirically assess the effects of data enrichment in various generative tasks, particularly in LRL contexts where original content is scarce but supplementary metadata may be available.
@@ 4.4. RQ4: Which Sources Are Publishing in the Area of Low-Resource Language Modelling?
While the results section for RQ4 indicates a wide variety of sources publishing in the area of LRL modelling, the sources themselves serve different purposes. Given the dominance of ACL publications spanning journals, workshops and conferences alike, it is clear that it has a central influence on the dissemination of literature relating to modelling LRLs. The high frequency of ACL publications from specific events such as the Empirical Methods in Natural Language Processing (EMNLP) and the Language Resources and Evaluation Conference (LREC) could suggest that these are key venues for LRL research. The prevalence of EMNLP papers could be attributed to the empirical nature of language modelling, and the frequency of LREC papers could be linked to the language resource element of building language models for LRLs. While ACM is associated with fewer events in comparison with other organisations such as ACL and IEEE, the Transactions on Asian and Low-Resource Language Information Processing journal is the most frequently sourced publication in this review, which could highlight that this journal is highly influential in terms of LRL modelling.
While arXiv preprint papers are not necessarily associated with traditional academic journals or undergo formal peer review, their prevalence could suggest that the field of LRL modelling is a rapidly evolving and current research topic. However, the frequency of arXiv papers could introduce limitations relating to the quality of research in the review process. We mitigate this potential bias by using the quality assessment questions relating to the rankings of conferences and journals. On the other hand, IEEE has more of a multi-disciplinary influence as it is the source of eight different publications relating to various fields within ICT and engineering. This diversity of publications could highlight that there are potential practical applications for LRL modelling in a variety of domains beyond purely theoretical research. Considering that Springer is the source of five journal papers in this review, this could indicate that there is a demand for topic-specific journals that relate to tackling LRL modelling. Furthermore, individual contributions from other publishers such as Frontiers, the International Joint Conferences on Artificial Intelligence, and De Gruyter may suggest that the research area of LRL modelling is gaining broader recognition and rising in prevalence across publishers. Ultimately, the broad range of sources that span eight different publishers indicates that there is strong academic interest in the field of LRL modelling.
@@ 4.5. RQ5: What Architectures Are Being Used to Produce Low-Resource Language Models?
The prevalence of transformer-based models in this review reflects the transformative nature of the architecture for generative language modelling. Given the rise in popularity, adoption and performance of transformer-powered services such as ChatGPT, it is unsurprising that it dominates as the most frequently implemented model architecture. The sheer popularity of transformer-based models suggests that transformers are the most suitable architecture for modelling LRLs. The significant drop in popularity for RNNs, and more specifically LSTMs could reflect the evolution of the state-of-the-art architecture over time as RNN/LSTM-based approaches would have been the gold standard before the transformer was introduced (Mienye et al., 2024). While relatively unpopular methods GAN, GRU and SMT-based methods have been used by studies in this review, their unpopularity could suggest that they are not the most optimal architectures for LRL generation in contrast with transformers. The limited use of these methods could reflect some attempts to use alternatives to the compute-heavy transformer architecture. For LRL communities that do not have access to sufficient computational power, such methods may still hold value as lightweight alternatives.
In terms of specific models, the popularity of GPT-2, the open-source predecessor to the widely popular GPT-3 and GPT-4 models by OpenAI, could indicate that it is suitable for LRL modelling. The relatively small size of GPT-2 small could make it appealing to LRL researchers, as it does not require as much training time and computational power to build it as other prevalent models such as the Llama3 family of models (Radford et al., 2019; Dubey et al., 2024). The frequent use of XLM and sequence-to-sequence models such as mT5 and T5 could be attributed to the considerable number of studies in this review that tackle translation, an inherently sequence-to-sequence task. Ultimately, while there is a wide variety of models being used to build generative models for LRLs, transformer-based architectures dominate as the most frequently used for this task. Future work could focus on optimising transformer architectures for LRL modelling tasks. Given the potential lack of computational resources in LRL communities, investigating model efficiency improvements could lower the barrier to entry for research in the area.
@@
While some text data augmentation techniques are effective at improving the performance of different NLG models, it is worth considering how such methods compare across different NLG tasks. This review finds that translation as a generative task can be improved with several augmentation approaches such as back-translation (Pham et al., 2021; Hong et al., 2024; Berckmann and Hiziroglu, 2020; Yirmibeşoğlu and Güngör, 2023; Nissanka et al., 2020; Dandapat and Federmann, 2018; Pham Van and Le Thanh, 2022), code-mixing (Bhowmick et al., 2023) and paraphrasing (Marie and Fujita, 2020; Guo et al., 2021a). It is unlikely that back-translation would be an effective method for bolstering the amount of training data for building models for other LRL NLG tasks such as summarisation or dialogue generation. Using back-translation for data synthesis could introduce semantic drift and inconsistencies at a syntactic level in the training corpus. Given that synthesising code mixed data has yielded performance increases for translation tasks, it could be insightful to apply this approach to languages that are often interwoven with each other. For example, some forms of modern Gaeilge usage could be more accurately modelled if it was trained on a corpus code mixed to include English, given the strong relationship between both languages (Laoire, 2016). Therefore, it could be insightful to investigate the impact of code-mixing on other NLG tasks. While paraphrasing has produced performance improvements for translation tasks, it has also yielded improvements for other tasks such as question-answering corpus generation and dialogue generation (Sun et al., 2019; Gao et al., 2020; Wu and Wang, 2023). This approach of using paraphrasing models to bolster training data is a promising approach that hinges on the availability of enough data or existing paraphrasing models. More research and experimentation are required to assess the impact of large-scale paraphrasing as a means to augment training data for building NLG models for LRLs. Training multilingual and language family-based models in particular has been quite effective at bridging the data gap between LRLs. Given that this language family-based approach can reduce the number of model parameters (Cahyawijaya et al., 2021) and that multilingual modelling can be compared to a 33% increase in training data for an LRL (Chang et al., 2023), it is clear that this is an effective text data augmentation method. It would be worth considering applying this multilingual, language family-based approach to various groups of related LRLs to further validate this observation. However, future work would be required to identify maximally related language family groupings and to collate existing resources across linguistic communities. While this multilingual, language family-based approach is promising from a general language modelling perspective, more research is required to assess its impact on generative task-specific performance.